The average number of cars arriving at a particular red light each day is 4. Assuming a Poisson distribution, calculate the probability that on a given day, less than 3 cars will arrive at the red light.


1
Expert's answer
2021-11-18T18:27:54-0500

By condition, "\lambda = 4" .

Let X represent the number of such cars.

Then

"P(X = 0) = \frac{{{\lambda ^0}}}{{0!}}{e^{ - \lambda }} = {e^{ - 4}}"

"P(X = 1) = \frac{{{\lambda ^1}}}{{1!}}{e^{ - \lambda }} = 4{e^{ - 4}}"

"P(X = 2) = \frac{{{\lambda ^2}}}{{2!}}{e^{ - \lambda }} = 8{e^{ - 4}}"

Then the wanted probability is

"P(X < 3) = P(X = 0) + P(X = 1) + P(X = 2) = 13{e^{ - 4}} \approx 0.238"

Answer: "P(X < 3) \approx 0.238"